OpenStudy (anonymous):

Verify the following for n=1 & n=3 1+5+9+...+(4n-3)=n(2n-1) and prove it where n belongs to Natural numbers by mathematical induction

OpenStudy (anonymous):

Essentially show that 1 = n(2n-1) where n=1 and 1+5+9=n(2n-1) where n=3

OpenStudy (anonymous):

prove it by induction for k+1

OpenStudy (anonymous):

1+5+9+...+(4k-3)=k(2k-1)

OpenStudy (anonymous):

add 4k+1 on both sides

OpenStudy (anonymous):

=k(2k-1)+4k+1

OpenStudy (anonymous):

2k^2-k+4k+1 <--- stuck here

OpenStudy (anonymous):

@Xavier

OpenStudy (anonymous):

You don't need to do anything complicated. Evaluate n(2n-1) at n=3 and show that equals 1+5+9

OpenStudy (anonymous):

sorry but I didn't mentioned the full question

OpenStudy (anonymous):

@Xavier ??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

hello anyone there??

OpenStudy (anonymous):

Ah sorry. General proof amounts to showing n(2n-1)+(4(n+1)-3) is the same as (n+1)(2(n+1)-1)

OpenStudy (anonymous):

yes I know but I am stuck

OpenStudy (anonymous):

Just foil each of them out

OpenStudy (anonymous):

what

OpenStudy (anonymous):

Do you see how I got the two equations I listed there?

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

Then expand them

OpenStudy (anonymous):

but I have to prove reverse way

OpenStudy (anonymous):

1+5+9+...+(4n-3)=n(2n-1)

OpenStudy (anonymous):

(4n-3) (4(n+1)-3) 4n+1 Add 4n+1 on bs

OpenStudy (anonymous):

1+5+9+...+(4n-3)+(4n+1)=n(2n-1)+(4n+1)

OpenStudy (anonymous):

2n^2-n+4k+1 <---- stuck here................ . . we have to make it (n+1)[2(n+1)-1]

OpenStudy (anonymous):

This is how I think about induction. Prove the equation holds for base case: n=1 Assume it holds for n: 1+5+9+...+(4n-3)=n(2n-1) is true Show it holds for n+1. 1+5+9+...+(4n-3) + (4n+1)=(n+1)(2(n+1)-1) See I substituted the n on the right with n+1. Now I know what 1+5+9+...+(4n-3) on the left is. n(2n-1)+(4n+1) = (n+1)(2(n+1)-1). Now is this true?